# Large Language Models in Argument Mining: A Survey

### v5

@@ 4 Dataset and Benchmark
ArgMinNews 2015 80 DE Text Argument Relation Identification
UKP OpposingArgumentsInEssays 2016 402 EN Text Claim Detection
UKP ConvArg 2016 16k EN Text Convincingness Checking
UKP InsufficientArguments 2017 1 k EN Text Sufficiency Checking
IBM Claim Stance Dataset 2017 2.3k EN Text Stance Detection
IBM Claim Sentences Search 2018 1.49M EN Text Claim Detection
IBM Evidence Quality 2019 5.6K EN Text Claim Detection
PERSPECTRUM 2019 20k EN Text Claim Detection & Evidence Detection
UKP ASPECT 2019 3.5K EN Text Claim Detection
IBM Labeled Emphasized Words in Speech 2018 2.4k EN Multi Text to Speech
AMPERSAND 2020 2.7k EN Text Argument Relation Identification
CoPA 2019 800 EN Text Claim Detection
IBM ArgQ-14kArgs 2019 14k EN Text Convincingness Checking
IBM Argumentative Sentences in Recorded Debates 2020 700 EN Text Claim Detection
Review-Rebuttal 2020 4.7K EN Text Claim Detection
IBM Evidences Sentences 2020 29k EN Text Convincingness Checking
IBM XArgMining 2020 35k DE,NL,E S,FR,IT Text Claim Detection & Evidence Detection & Stance Detection
IBM Recorded Debating Dataset 2020 3.5k EN Multi Speech to Text
MEDLINE 2020 6.7k EN Text Claim Detection & Argument Relation Identification
GAQCorpus 2020 5k EN Text Convincingness Checking & Sufficiency Checking
IBM ArgKP 2021 27k EN Text Key Point Analysis
SciARK 2021 1k EN Text Claim Detection & Evidence Detection
IAM 2022 69k EN Text Claim Detection & Evidence Detection & Stance Detection
FinArg 2022 136 EN Text Claim Detection & Stance Detection
Sustainable Diet Arguments on Twitter 2022 592 EN Text Claim Detection & Evidence Detection & Stance Detection
@@
Kialo 2022 310k EN Text Argument Relation Identification
IBM SurveyKP 2023 15k EN Text Key Point Analysis
ArguGPT 2023a 8k EN Text Claim Detection
StoryARG 2023 2.4k EN Text Claim Detection & Argument Relation Identification
GENEVA 2023 3k EN Text Argument Relation Identification
VivesDebate-Speech 2023 29 EN Multi Speech to Text
ArgSciChat 2023 498 EN Text Argument Relation Identification
MuLMS-AZ 2023 10k EN Text Argument Relation Identification
ArgAnalysis35K 2023 35k EN Text Argument Relation Identification
ORCHID 2023 14k EN Text Argument Summarization & Stance Detection
ARIES 2024 81k EN Text Argument Relation Identification
FS150T Corpus 2024 21k EN Text Claim Detection & Stance Detection
CasiMedicos-Arg 2024 7k EN Text Claim Detection & Stance Detection
CEAMC 2024 226 EN,ZH Text Claim Detection & Evidence Detection & Stance Detection
ASE 2024b 2.3 k EN Text Evidence Detection & Convincingness Checking & Argument Summarization & Sufficiency Checking
ConQRet 2025 6.5k EN Text Evidence Detection & Argument Quality Assessment
@@
Specifically, we outline it in three successive research phases: (i) from task-specific, human-curated collections to integrated, LLM-centric resource (Section 4.1), (ii) Using LLM as annotator and generator (Section 4.2, and (iii) an emerging focus on scale, modality, and language breadth (Section 4.3). We review each phase and close look at over 40 datasets (shown in Table 1 and 2 to figure out the trends and open challenges.
@@ 4.1 From One Task–One Dataset to Integrated Corpora
[…]
@@
Single-purpose design proved invaluable for methodological clarity. Because each dataset held only one variable constant, researchers could dissect feature contributions and model architectures with minimal confounds. Notably, margin-ranking SVMs on ConvArg delivered a 19-point Pairwise Accuracy gain over bag-of-words in large part because the dataset’s simplicity isolated ranking logic from component detection errors (Habernal and Gurevych, 2016). However, this fragmentation soon impeded cumulative progress. Cross-task transfer became almost impossible: a model trained to find claims in UKP essays failed on biomedical abstracts, not because the notion of a claim differed but because no shared label inventory or evaluation split existed across corpora (Mayer et al., 2020). Moreover, meta-analysis was thwarted by incompatible annotation units (sentence vs. clause) and divergent definitions of “support” and “attack” (Reimers et al., 2019).
@@
Faced with these limitations—and armed with instruction-tuned large language models capable of multi-task learning—dataset designers began to integrate multiple label layers within a single corpus. The turning point was IAM (69 k Reddit threads), which simultaneously marks clause boundaries, supporting evidence links, stance polarity, and topic metadata. Multi-task RoBERTa trained on IAM improved macro-F1 by 4–6 points over single-task baselines on both claim detection and stance classification, demonstrating that shared representations benefit from joint supervision (Cheng et al., 2022). Similarly, ARIES (81 k documents) overlays RST-style discourse graphs on argumentative zones; graph neural networks trained here transfer to the cross-lingual MuLMS-AZ news corpus with only a 5-point F1 loss, far less than the 15-point drops reported for earlier single-task systems (Gemechu et al., 2024; Schrader et al., 2023). Integrated corpora are not confined to component-level annotation: ORCHID fuses stance, key-point matching, and abstractive summaries in 14 k opinion paragraphs, enabling the first evaluation of summarisation metrics grounded in argumentative structure (Zhao et al., 2023); StoryARG does likewise for narrative argumentation (Falk and Lapesa, 2023).
[…]
@@ 4.2 LLM-Assisted Creation and Annotation
gained prominence when Liu et al. (2023a) used GPT-4 to draft thousands of claim–counter-claim pairs across diverse topics. Although human editors rejected a non-trivial minority for logical incoherence, the retained material exhibits argumentative structures broad enough to boost zero-shot performance on earlier convincingness benchmarks by a noticeable margin—often up to ten percent relative improvement. Stylistic audits, however, reveal a homogenised, over-polite register and a marked reduction in hedging compared with organically written debate, raising concerns that synthetic data may narrow the stylistic bandwidth of subsequent models (Long et al., 2024).
@@
Humans remain in control but delegate first-pass labelling to an LLM. The ASE corpus exemplifies this workflow: GPT-4 proposes evidence spans, sufficiency judgements and short summaries, after which human reviewers confirm or amend the suggestions (Li et al., 2024b). Verification time falls markedly, yet error analysis shows that chain-of-thought prompting, while improving recall, doubles the rate of subtle hallucinations. Similar hybrid gains are reported for SurveyKP, where key-point clusters refined by annotators achieve noticeably higher purity than crowd-only baselines at roughly half the labour hours (Eden et al., 2023). The flip side is automation anchoring: reviewers tend to accept stylistic quirks or soft politeness biases introduced by the model, an effect already quantified in appropriateness studies by Ziegenbein et al. (2023).
@@
A third trend outsources to automatic judgements LLMs when labels are cognitively heavy or prone to low consensus. In ConQRet, GPT-4 decides whether retrieved passages “fully answer” complex questions before experts handle only borderline cases (Dhole et al., 2025). Agreement with specialists is high overall, yet topic-sensitive analyses show the model penalises rhetorical questions more harshly than humans—a bias echoing manipulation-check findings on emotional language by Chen and Eger (2025). Moreover, when the same model family later evaluates systems trained on these labels, circularity becomes a real possibility: sufficiency scores judged by GPT-4 appear several points higher when the evaluator is also GPT-4 than when a conventional RoBERTa baseline is used (Liu et al., 2024b).
@@ 4.3 Scale, Modality, and Language
The median pre-LLM corpus contained only a few thousand annotated instances, limiting the depth of neural fine-tuning. Integrated resources now appear at an order-of-magnitude larger scale. Debate platform dumps such as Kialo  [^5] https://www.kialo.com/, with well over a quarter-million crowd-moderated nodes, allow graph-level experiments in argumentation theory that were inconceivable when researchers relied on essay corpora (Agarwal et al., 2022). Large scale is not merely a matter of quantity: it supports adversarial splitting, more granular error analysis, and realistic low-resource simulations. Yet volume can disguise annotation drift; agreement figures reported for the first 10
@@
Argumentative discourse is rarely confined to prose; tone of voice, slide layouts, or imagery all shape persuasiveness. Two speech–text corpora IBM’s Recorded Debating and the VivesDebate-Speech set align transcribed turns with audio and meta-data, enabling first forays into prosody-aware argument strength (Orbach et al., 2020; Ruiz-Dolz and Sanchez, 2023). The Image-Argument pilot goes further, annotating visual claims and layout cues in news pages (Liu et al., 2023b). Early multimodal baselines show detectable gains when acoustic or layout features are fused with text, but they also reveal new annotation pain-points: irony carried by image–caption incongruence or persuasive pauses in speech remain largely uncaptured. More importantly, evaluation metrics such as “pairwise convincingness” have yet to be re-interpreted for cross-modal signals.
@@
English continues to dominate, but a second tier of languages,German, Dutch, Spanish, French, Italian, and Chinese, has emerged. Parallel corpora such as XArgMining demonstrate that simple cross-lingual projection recovers much, but not all, of English performance (Toledo-Ronen et al., 2020). Specialised bilingual resources add depth: FinArg pairs English and domain-specific finance jargon with argumentative structure (Alhamzeh et al., 2022), whereas CEAMC offers Chinese–English medical arguments linked by explicit claim–evidence pairs (Ren et al., 2024). Nevertheless, most multi-task corpora remain monolingual, so joint modelling of component, stance and quality across scripts is still aspirational. Preliminary cross-lingual experiments suggest that stance and sufficiency transfer degrade much faster than claim segmentation, hinting that cultural framing affects higher-order argumentative facets more than low-level span boundaries.
@@
The three expansion axes seldom intersect in the same dataset: no existing corpus is simultaneously large-scale, multimodal, and multilingual. This fragmentation hampers holistic robustness tests—for instance, a model trained on English speech debates cannot yet be evaluated on Chinese medical videos. Future benchmarks should therefore prioritise cross-dimensional design, even if that means curating smaller but more diverse seed sets. Equally urgent is the development of evaluation metrics that remain meaningful when text, audio, and visuals interact. Without such metrics, gains reported on one modality may mask losses in another, leading to brittle real-world systems.

### v6

@@ 3 Dataset and Benchmark
ArgMinNews DBLP:conf/emnlp/Eckle-KohlerKG15 80 DE Text Argument Relation Identification
UKP OpposingArgumentsInEssays DBLP:conf/acl/StabG16 402 EN Text Claim Detection
UKP ConvArg DBLP:conf/acl/HabernalG16 16k EN Text Convincingness Checking
UKP InsufficientArguments DBLP:conf/eacl/GurevychS17 1 k EN Text Sufficiency Checking
IBM Claim Stance Dataset DBLP:conf/eacl/SlonimBSBD17 2.3k EN Text Stance Detection
IBM Claim Sentences Search DBLP:conf/coling/LevyBGAS18 1.49M EN Text Claim Detection
IBM Evidence Quality DBLP:conf/acl/GleizeSCDMAS19 5.6K EN Text Claim Detection
PERSPECTRUM DBLP:conf/naacl/ChenK0CR19 20k EN Text Claim Detection & Evidence Detection
UKP ASPECT DBLP:conf/acl/ReimersSBDSG19 3.5K EN Text Claim Detection
IBM Labeled Emphasized Words in Speech DBLP:conf/interspeech/MassSMHSLK18 2.4k EN Multi Text to Speech
AMPERSAND DBLP:journals/corr/abs-2004-14677 2.7k EN Text Argument Relation Identification
CoPA DBLP:conf/acl/BiluGHSLMMGS19 800 EN Text Claim Detection
IBM ArgQ-14kArgs DBLP:conf/emnlp/ToledoGCFVLJAS19 14k EN Text Convincingness Checking
IBM Argumentative Sentences in Recorded Debates DBLP:conf/emnlp/ShnarchCMAS20 700 EN Text Claim Detection
Review-Rebuttal DBLP:conf/emnlp/ChengBYLS20 4.7K EN Text Claim Detection
IBM Evidences Sentences DBLP:conf/aaai/Ein-DorSDHSGAGC20 29k EN Text Convincingness Checking
IBM XArgMining DBLP:conf/emnlp/Toledo-RonenOBS20 35k DE,NL,E S,FR,IT Text Claim Detection & Evidence Detection & Stance Detection
IBM Recorded Debating Dataset DBLP:conf/acl/OrbachBTLJAS20 3.5k EN Multi Speech to Text
MEDLINE DBLP:conf/ecai/0002CV20 6.7k EN Text Claim Detection & Argument Relation Identification
GAQCorpus DBLP:conf/coling/LauscherNNT20 5k EN Text Convincingness Checking & Sufficiency Checking
IBM ArgKP DBLP:conf/argmining/FriedmanDHAKS21 27k EN Text Key Point Analysis
SciARK DBLP:conf/argmining/FergadisPKP21 1k EN Text Claim Detection & Evidence Detection
IAM DBLP:conf/acl/ChengBHYZS22 69k EN Text Claim Detection & Evidence Detection & Stance Detection
FinArg alhamzeh2022s 136 EN Text Claim Detection & Stance Detection
Sustainable Diet Arguments on Twitter hansen2022dataset 592 EN Text Claim Detection & Evidence Detection & Stance Detection
@@
Kialo agarwal2022graphnli 310k EN Text Argument Relation Identification
IBM SurveyKP DBLP:conf/emnlp/EdenKOKSB23 15k EN Text Key Point Analysis
ArguGPT DBLP:journals/corr/abs-2304-07666 8k EN Text Claim Detection
StoryARG DBLP:conf/acl/FalkL23 2.4k EN Text Claim Detection & Argument Relation Identification
GENEVA DBLP:conf/acl/ParekhHHCP23 3k EN Text Argument Relation Identification
VivesDebate-Speech DBLP:conf/emnlp/Ruiz-DolzS23 29 EN Multi Speech to Text
ArgSciChat DBLP:conf/acl/RuggeriMG23 498 EN Text Argument Relation Identification
MuLMS-AZ DBLP:journals/corr/abs-2307-02340 10k EN Text Argument Relation Identification
ArgAnalysis35K DBLP:conf/acl/JoshiPH23 35k EN Text Argument Relation Identification
ORCHID DBLP:conf/emnlp/ZhaoWP23 14k EN Text Argument Summarization & Stance Detection
ARIES gemechu2024aries 81k EN Text Argument Relation Identification
FS150T Corpus DBLP:conf/emnlp/SchillerDWG24 21k EN Text Claim Detection & Stance Detection
CasiMedicos-Arg DBLP:conf/emnlp/SviridovaYECVA24 7k EN Text Claim Detection & Stance Detection
CEAMC DBLP:conf/emnlp/RenWLZZYZBL24 226 EN,ZH Text Claim Detection & Evidence Detection & Stance Detection
ASE DBLP:conf/acl/0074WSBMZZWHLN24 2.3 k EN Text Evidence Detection & Convincingness Checking & Argument Summarization & Sufficiency Checking
ConQRet DBLP:conf/naacl/DholeSA25 6.5k EN Text Evidence Detection & Argument Quality Assessment
@@
Specifically, we outline it in three successive research phases: (i) from task-specific, human‐curated collections to integrated, LLM-centric resource (Section 3.1), (ii) Using LLM as annotator and generator (Section 3.2, and (iii) an emerging focus on scale, modality, and language breadth (Section 3.3). We review each phase and close look at over 40 datasets (shown in Table 1 and 2 to figure out the trends and open challenges.
@@ 3.1 From One Task–One Dataset to Integrated Corpora
The first generation of argument-mining resources was intentionally mono-task. Each corpus addressed a single research question and encoded only the labels required to answer it. For instance, UKP OpposingArgumentsInEssays focussed exclusively on segmenting major claims and premises in student essays and reached expert agreement of $\kappa=.81$ by concentrating annotator effort on two labels alone (DBLP:conf/acl/StabG16). UKP ConvArg framed argument quality as pairwise convincingness comparisons, enabling early feature-based rankers and later RoBERTa baselines, but left logical sufficiency and rhetorical style unlabelled (DBLP:conf/acl/HabernalG16). IBM InsufficientArguments captured binary sufficiency in 1 000 essay sentences; the narrow scope made it ideal for entailment studies, yet its labels could not generalise to social-media debates or legal opinions (DBLP:conf/eacl/GurevychS17).
@@
Single-purpose design proved invaluable for methodological clarity. Because each dataset held only one variable constant, researchers could dissect feature contributions and model architectures with minimal confounds. Notably, margin-ranking SVMs on ConvArg delivered a 19-point Pairwise Accuracy gain over bag-of-words in large part because the dataset’s simplicity isolated ranking logic from component detection errors (DBLP:conf/acl/HabernalG16). However, this fragmentation soon impeded cumulative progress. Cross-task transfer became almost impossible: a model trained to find claims in UKP essays failed on biomedical abstracts, not because the notion of a claim differed but because no shared label inventory or evaluation split existed across corpora (DBLP:conf/ecai/0002CV20). Moreover, meta-analysis was thwarted by incompatible annotation units (sentence vs. clause) and divergent definitions of “support” and “attack” (DBLP:conf/acl/ReimersSBDSG19).
@@
Faced with these limitations—and armed with instruction-tuned large language models capable of multi-task learning—dataset designers began to integrate multiple label layers within a single corpus. The turning point was IAM (69 k Reddit threads), which simultaneously marks clause boundaries, supporting evidence links, stance polarity, and topic metadata. Multi-task RoBERTa trained on IAM improved macro-F1 by 4–6 points over single-task baselines on both claim detection and stance classification, demonstrating that shared representations benefit from joint supervision (DBLP:conf/acl/ChengBHYZS22). Similarly, ARIES (81 k documents) overlays RST-style discourse graphs on argumentative zones; graph neural networks trained here transfer to the cross-lingual MuLMS-AZ news corpus with only a 5-point F1 loss, far less than the 15-point drops reported for earlier single-task systems (gemechu2024aries; DBLP:journals/corr/abs-2307-02340). Integrated corpora are not confined to component-level annotation: ORCHID fuses stance, key-point matching, and abstractive summaries in 14 k opinion paragraphs, enabling the first evaluation of summarisation metrics grounded in argumentative structure (DBLP:conf/emnlp/ZhaoWP23); StoryARG does likewise for narrative argumentation (DBLP:conf/acl/FalkL23).
The move to multi-facet datasets raises new methodological questions. Annotator agreement is now label-specific: in IAM, component spans reach $\alpha=.79$, yet stance falls to .63, and evidence links to .57—numbers rarely reported in earlier corpora. Label interoperability remains ad hoc: the definition of “premise” in IAM (full clause) does not exactly match that in ARIES (minimal span), complicating model reuse. Preliminary work on ontology mapping (DBLP:conf/emnlp/IvanovaHN24) shows that automatic label harmonisation lifts zero-shot transfer by 6–8 points, but the field still lacks a canonical schema.
@@ 3.2 LLM-Assisted Creation and Annotation
gained prominence when DBLP:journals/corr/abs-2304-07666 used GPT-4 to draft thousands of claim–counter-claim pairs across diverse topics. Although human editors rejected a non-trivial minority for logical incoherence, the retained material exhibits argumentative structures broad enough to boost zero-shot performance on earlier convincingness benchmarks by a noticeable margin—often up to ten percent relative improvement. Stylistic audits, however, reveal a homogenised, over-polite register and a marked reduction in hedging compared with organically written debate, raising concerns that synthetic data may narrow the stylistic bandwidth of subsequent models (long2024llms).
@@
Humans remain in control but delegate first-pass labelling to an LLM. The ASE corpus exemplifies this workflow: GPT-4 proposes evidence spans, sufficiency judgements and short summaries, after which human reviewers confirm or amend the suggestions (DBLP:conf/acl/0074WSBMZZWHLN24). Verification time falls markedly, yet error analysis shows that chain-of-thought prompting, while improving recall, doubles the rate of subtle hallucinations. Similar hybrid gains are reported for SurveyKP, where key-point clusters refined by annotators achieve noticeably higher purity than crowd-only baselines at roughly half the labour hours (DBLP:conf/emnlp/EdenKOKSB23). The flip side is automation anchoring: reviewers tend to accept stylistic quirks or soft politeness biases introduced by the model, an effect already quantified in appropriateness studies by DBLP:conf/acl/ZiegenbeinSLPW23.
@@
A third trend outsources to automatic judgements LLMs when labels are cognitively heavy or prone to low consensus. In ConQRet, GPT-4 decides whether retrieved passages “fully answer” complex questions before experts handle only borderline cases (DBLP:conf/naacl/DholeSA25). Agreement with specialists is high overall, yet topic-sensitive analyses show the model penalises rhetorical questions more harshly than humans—a bias echoing manipulation-check findings on emotional language by DBLP:journals/corr/abs-2503-00024. Moreover, when the same model family later evaluates systems trained on these labels, circularity becomes a real possibility: sufficiency scores judged by GPT-4 appear several points higher when the evaluator is also GPT-4 than when a conventional RoBERTa baseline is used (DBLP:conf/naacl/LiuFC24).
@@ 3.3 Scale, Modality, and Language
The median pre-LLM corpus contained only a few thousand annotated instances, limiting the depth of neural fine-tuning. Integrated resources now appear at an order‐of-magnitude larger scale. Debate platform dumps such as Kialo  [^5] https://www.kialo.com/, with well over a quarter-million crowd-moderated nodes, allow graph-level experiments in argumentation theory that were inconceivable when researchers relied on essay corpora (agarwal2022graphnli). Large scale is not merely a matter of quantity: it supports adversarial splitting, more granular error analysis, and realistic low-resource simulations. Yet volume can disguise annotation drift; agreement figures reported for the first 10
@@
Argumentative discourse is rarely confined to prose; tone of voice, slide layouts, or imagery all shape persuasiveness. Two speech–text corpora IBM’s Recorded Debating and the VivesDebate-Speech set align transcribed turns with audio and meta-data, enabling first forays into prosody-aware argument strength (DBLP:conf/acl/OrbachBTLJAS20; DBLP:conf/emnlp/Ruiz-DolzS23). The Image-Argument pilot goes further, annotating visual claims and layout cues in news pages (DBLP:conf/argmining/LiuEZL23). Early multimodal baselines show detectable gains when acoustic or layout features are fused with text, but they also reveal new annotation pain-points: irony carried by image–caption incongruence or persuasive pauses in speech remain largely uncaptured. More importantly, evaluation metrics such as “pairwise convincingness” have yet to be re-interpreted for cross-modal signals.
@@
English continues to dominate, but a second tier of languages,German, Dutch, Spanish, French, Italian, and Chinese, has emerged. Parallel corpora such as XArgMining demonstrate that simple cross-lingual projection recovers much, but not all, of English performance (DBLP:conf/emnlp/Toledo-RonenOBS20). Specialised bilingual resources add depth: FinArg pairs English and domain-specific finance jargon with argumentative structure (alhamzeh2022s), whereas CEAMC offers Chinese–English medical arguments linked by explicit claim–evidence pairs (DBLP:conf/emnlp/RenWLZZYZBL24). Nevertheless, most multi-task corpora remain monolingual, so joint modelling of component, stance and quality across scripts is still aspirational. Preliminary cross-lingual experiments suggest that stance and sufficiency transfer degrade much faster than claim segmentation, hinting that cultural framing affects higher-order argumentative facets more than low-level span boundaries.
@@
These expansion axes rarely intersect: no existing corpus is simultaneously large-scale, multimodal, and multilingual. This limits robustness testing—for example, a model trained on English audio debates cannot be evaluated on Chinese medical videos. Future benchmarks should therefore pursue cross-dimensional design, even at reduced scale. Equally urgent is the development of evaluation metrics that remain meaningful when text, audio, and images interact; without such metrics, improvements in one modality may conceal failures in another.